OpenStudy (anonymous):

Solve the system, using substitution. 3a - b = 11 and 2a + 3b = 0 A. (3, 2) B. (3, -2) C. (-3, 2) D. (-3, -2) E. no solution

OpenStudy (anonymous):

b=3a-11 2a+3(3a-11)=0 2a+9a-33=0 11a=33 a=3 2(3)+3b=0 6+3b=0 b=2 A is your answer

OpenStudy (texaschic101):

b = 3a - 11 2a + 3(3a - 11) = 0 2a + 9a - 33 = 0 11a - 33 = 0 11a = 33 a = 3 2(3) + 3b = 0 6 + 3b = 0 3b = -6 b = -2 answer is (3,-2)....answer is B

OpenStudy (anonymous):

You're right. Thanks for pointing that out. I'm going too fast x.x

OpenStudy (texaschic101):

lol....no problem...if I had $1 for every mistake I made, I would be rich